OpenStudy (anonymous):

For the polar curve r=3cos(2theta) find the locations of all horizontal and vertical tangents

OpenStudy (anonymous):

The issue was that this one really blew up in my face when I found dy/dx, and am having a hard time finding where it equals zero/is undefined

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

r=3cos(2t) rr=3 rcos(2t) x^2 + y^2 =3 r(cos^2(t)-sin^2(t)) x^2 + y^2 =3 (rcos^2(t)-rsin^2(t)) x^2 + y^2 =3 (x cos(t)-y sin(t)) 2x x' + 2y y' =3 (x' cos(t)-xt' sin(t)-y' sin(t)-yt' cos(t)) -------------------------------------------- dy/dx..... let x=1, y' = dy/dx 2x + 2y y' =3cos(t) -3xt' sin(t) -3y' sin(t) -3yt' cos(t)) 3y' sin(t) + 2y y' =3cos(t) -3xt' sin(t) -3yt' cos(t)) - 2x y'(3sin(t) + 2y) =3cos(t) -3xt' sin(t) -3yt' cos(t)) - 2x y' =(3cos(t) -3xt' sin(t) -3yt' cos(t)) - 2x)/(3sin(t) + 2y) --------------------------------------------- when the top is zero we are tangent horizontal when the bottom is zero we are tangent vertical --------------------------------------------- bottom seems simplest 3sin(t) + 2y = 0 3sin(t) + 2r sin(t) = 0 3sin(t) + 6cos(2t)sin(t) = 0 3sin(t) (1+ 2cos(2t)) = 0 sin(t) = 0 or cos(2t) = -1/2 ------------------------------------------ top part, lets give it a whirl 3cos(t) -3xt' sin(t) -3yt' cos(t) - 2x = 0 3cos(t) -3rt' cos(t)sin(t) -3rt' sin(t) cos(t) - 2r cos(t) = 0 cos(t) (1 -6t' cos(2t) sin(t) - 2cos(2t)) = 0 cos(t) = 0 is one spot 1 -6t' cos(2t) sin(t) - 2cos(2t) = 0 this one might be trickiest :)

OpenStudy (anonymous):

You got a different dy/dx than I did

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=polar+plot+r%3D3cos%282t%29 the pictures nice :)

OpenStudy (anonymous):

I found dy/dtheta and dx/dtheta by using x=rcostheta and y=rsintheta. Then I substituted r into each equation and got x=(3cos(2theta))costheta and y=(3cos(2theta))sintheta

OpenStudy (anonymous):

Then I said dy/dx=(dy/dtheta)(dx/dtheta), and I derived my x and y equations. Which exploded...

OpenStudy (amistre64):

dy = dy dt dx dt dx

OpenStudy (anonymous):

Does t=theta?

OpenStudy (amistre64):

of course

OpenStudy (anonymous):

Oh yeah, I meant (dy/dtheta)OVER(dx/dtheta) :P Sorry

OpenStudy (anonymous):

But yeah, it got really complicated

OpenStudy (amistre64):

we can evaluate it from 0 to pi/4 and the rest of the graph is copies

OpenStudy (anonymous):

Evaluate what?

OpenStudy (amistre64):

the graph .... i posted what it looks like if you didnt take the time to sketch it

OpenStudy (anonymous):

Yeah, I know what it looks like. I have to prove the tangent lines though. For my derivative, I got dy/dx=\[\frac{ 3\cos(2\theta)\cos(\theta)+\sin(\theta)(-6\sin(2\theta)) }{ -3\cos(2\theta)\sin(\theta)-6\sin(2\theta)\cos(\theta)}\]

OpenStudy (anonymous):

Which looks different than what you got

OpenStudy (amistre64):

yeah i just went for the straight forward implicit derivative \[\frac{dy}{dx} =\frac{3cos(t) -3xt' sin(t) -3yt' cos(t) - 2x}{3sin(t) + 2y}\]

OpenStudy (anonymous):

Did we get the same thing though?

OpenStudy (amistre64):

dunno :) what do you get for your verticals?

OpenStudy (anonymous):

Thats the problem. I was trying to find the tangents and I didn't really know how to do that because of all of the trig functions

OpenStudy (amistre64):

when the slope is undefined, its vertical a fraction is undefined when its denom is zero 3sin(t) + 2y = 0, but y = r cos(t) = 3cos(2t) sin(t) 3sin(t) (1 + 2cos(2t)) = 0 its a product, so its zero when either of its factors or 0 sin(t) = 0, or cos(2t) = -1/2

OpenStudy (amistre64):

i cant verify that your derivative is good ....

OpenStudy (anonymous):

Can we use my derivative? That is the way that my teacher would rather see

OpenStudy (amistre64):

post all of your work to get there, and i can follow it

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

r=3cos(2theta)

OpenStudy (anonymous):

x=rcos(theta)=(3cos(2theta))cos(theta) y=rsin(theta)=(3cos(2theta))sin(theta)

OpenStudy (anonymous):

\[\frac{ dy }{ dx }=\frac{ 3\cos(2\theta)\cos(\theta)+\sin(\theta)(-3\sin(2\theta))2 }{ (3\cos(2\theta))(-\sin(\theta))+\cos(\theta)(-3\sin(2\theta))2 }\]

OpenStudy (anonymous):

And I simplified that a little bit.

OpenStudy (amistre64):

dy/dt = -6sin(2t) sin(t) + 3cos(2t) cos(t) = 0 give horizontal dx/dt = -6sin(2t)cos(t) - 3cos(2t)sin(t) = 0 give vertical agreed?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Exactly

OpenStudy (anonymous):

And I wasn't really sure how to get the zeros of those

OpenStudy (amistre64):

well dunno if itll simplfy it but cos(2t) = cos^2(t) - sin^2(t) =1 -2sin^2(t) =2 cos^2(t) - 1 sin(2t) = 2 sin(t)cos(t)

OpenStudy (anonymous):

I tried doing some stuff with those identities, but it just messed it up. I tried simplifying to get cot(2theta)=2tan(theta)

OpenStudy (anonymous):

But that doesn't really seem to work either

OpenStudy (amistre64):

-6sin(2t)cos(t) - 3cos(2t)sin(t) -12sin(t)cos(t)cos(t) - 3(2cos^2(2) - 1)sin(t) -12sin(t)cos^2(t) - 6sin(t)cos^2(t) +3sin(t) -18sin(t)cos^2(t) +3sin(t) 3sin(t)(-6cos^2(t) +1) = 0 sin(t) = 0, and 1/6 = cos^2(t); cos(t) = sqrt(6)/6 might be some errors in that along the way ....

OpenStudy (anonymous):

Well its not cos(theta)=root6/6

OpenStudy (anonymous):

oh wait. I forgot that the sides of the petals have tangents too

OpenStudy (anonymous):

I see! Thank you!!

OpenStudy (amistre64):

youre welcome

OpenStudy (amistre64):

and yes, cos(t) = rt6/6 is a zero on the bottom